OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 2) and is parallel to y = 4x + 7. A: y = 4x + 10 B: y= 1/4x+5/2 C: y = -4x - 6 D: y = -1/4x +3/2

OpenStudy (ybarrap):

Look for an equation with the same slope and plug in your coordinates to see if they match, you should get your answer very quickly like this

OpenStudy (ybarrap):

note that if lines are parallel, they will have the same slope. Think about it

OpenStudy (anonymous):

I still don't really understand it. I'm terrible with graphs and slopes. My best guess is that it is A. Cause it would be parallel but just to the side, right?

OpenStudy (texaschic101):

do you know what the slope is in the original equation ?

OpenStudy (ybarrap):

y=mx+b, m = slope, b=y-intercept (i.e. where the line crosses the y-axis when x=0)

OpenStudy (texaschic101):

you could just use point slope formula and sub in slope (m) and your points given

OpenStudy (anonymous):

I got C... But I'm very unsure about that one.

OpenStudy (ybarrap):



OpenStudy (anonymous):

I'm just going to guess D. Sorry, I guess i just dont get it

OpenStudy (texaschic101):

y = 4x + 7......what is your slope in this equation ?

OpenStudy (texaschic101):

ok...y = mx + b form, the slope is m. So in that equation , your slope is 4. A parallel line will have the same slope. Now we will use the point slope formula because we have the slope and a point. point slope formula is : y - y1 = m(x - x1) slope (m) = 4 (-2,2) x1 = -2, y1 = 2 now lets sub y - 2 = 4(x -(-2) y - 2 = 4(x + 2) -- distribute the 4 through the parenthesis y - 2 = 4x + 8 -- now add 2 to both sides to get y by itself y = 4x + 8 + 2 y = 4x + 10 <-- your answer Did that help you understand it better ?

OpenStudy (anonymous):

any line parallel to y=mx+c can be taken as y=mx+k so lets take the equation as y=4x+k now we have a point on the line (-2,2)lets put it into the equation 2=-8+k k=10 hence the equation is y = 4x +10 option a